148 Literaturverzeichnis

Literatur

[1] A. Einstein, Verh. Dt. Phys. Ges. 19, 82 (1917).

[2] L. Brillouin, J. Phys. Radium 7, 353 (1926).

[3] J. B. Keller, Ann. Phys. (N. Y.) 4, 180 (1958).

[4] V. P. Maslov, Theorie des Perturbations et Methodes Asmyptotiques (Dunod, Paris, 1972).

[5] J. H. Van Vleck, Proc. Natn. Acad. Sci. 14, 178 (1928).

[6] P. A. M. Dirac, The Principles of Quantum Mechanics (Clarendon Press, Oxford, 1947).

[7] M. V. Berry und K. E. Mount, Rep. Prog. Phys. 35, 315 (1972).

[8] R. P. Feynman und A. R. Hibbs, Quantum Mechanics and Path Integrals (Mc Graw–Hill, New York, 1965).

[9] L. S. Schulman, Techniques and Applications of Path Integration (Wiley, New York, 1981).

[10] M. C. Gutzwiller, J. Math. Phys. 12, 343 (1971).

[11] M. C. Gutzwiller, Chaos in Classical and Quantum Mechanics (Springer, New York, 1990).

[12] M. Tabor, Physica D 6, 195 (1983).

[13] G. Junker und H. Leschke, Physica D 56, 135 (1992).

[14] R. Balian und C. Bloch, Ann. Phys. (N. Y.) 69, 76 (1972).

[15] M. V. Berry und M. Tabor, Proc. R. Soc. Lond. A 349, 101 (1976).

[16] M. V. Berry und M. Tabor, J. Phys. A 10, 371 (1977).

[17] P. J. Richens, J. Phys. A 15, 2101 (1982).

[18] L. Markus und K. R. Meyer, Memoirs of the AMS 144 (American Mathematical Society, Providence, Rhode Island, 1974).

[19] F. Haake, M. Kuś, J. Mostowski und R. Scharf, in Coherence, Cooperation, and Fluctuations, herausgegeben von F. Haake, L. M. Narducci und D. F. Walls (Cambridge University Press, Cambridge, 1986), S. 220–230.

[20] F. Haake, M. Kuś und R. Scharf, Z. Phys. B 65, 381 (1987).

[21] F. Haake, Quantum Signatures of Chaos (Springer, Berlin, 1991).

[22] M. Kuś, F. Haake und B. Eckhardt, Z. Phys. B 92, 221 (1993).

[23] M. Kuś, F. Haake und D. Delande, Phys. Rev. Lett. 71, 2167 (1993).

[24] P. Gerwinski, F. Haake, M. Kuś, H. Wiedemann und K. Życzkowski, Phys. Rev. Lett. 74, 1562 (1995).

[25] P. A. Braun, P. Gerwinski, F. Haake und H. Schomerus, Z. Phys. B 100, 115 (1996).

[26] H. Schomerus und M. Sieber, J. Phys. A 30, 4537 (1997).

[27] M. Sieber und H. Schomerus, chao-dyn/9708013 (1997).

[28] H. Schomerus, Europhys. Lett. 38, 423 (1997).

[29] H. Schomerus und F. Haake, Phys. Rev. Lett. 79, 1022 (1997).